How would you make a 1:1000 dilution of E. coli?
Answer #1

Add 1ml of culture to 999 ml of water. You will get 1:1000 dilution.

Or add 1µl of cultue to 999 µl of water. You will get 1:1000 dilution.

D= 1/(1+999)=1/1000=1 x 10-3 (this is usually written as 1:1000 dilution)